A Complete Bibliography of Computability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

30 March 2021 Version 1.04

Title word cross-reference

+ [Mil17]. 0 ∼′ [Ger20]. < [Mil17]. ^{1} [HKS17]. ^{ℓ}p [McN17]. _{0} [SS15, dBSS16]. C[0, 1] [Nem20]. ℓp [MS19]. H [GJM^{+}19]. L^{1} [Miy13]. ω [Ger20]. ω_{1} [CJK^{+}13]. Q [Sel20]. T [Nor16].

-colouring [GJM^{+}19]. -Computability [Miy13]. -definable [Mil17]. -REA [Ger20]. -spaces [SS15]. -spaces1 [dBSS16]. -Wadge [Sel20].

2018 [BDKL19, BDMP20].

above [CDHTM20]. admissible [Joh20]. affine [BA15]. Algebraic [DHS13]. algebras [AZ19]. algorithm [CD18, CD20]. Algorithmic [FrKHNS14, STZDG13, Muc16]. algorithmically [HTKT19]. Analog [PZ18, Mil20]. analysis [San18]. anonymous [GL17]. anytime [CD18, CD20]. approach [Mil20]. approximable [KN21]. Approximation [Mor12]. arithmetic [Spe15]. Arithmetical [CJK^{+}13]. aspect [San18]. Aspects [FrKHNS14, HTKT19, Pau16]. Assembly [SO13]. Asymptotic [Ast15, HJMS16]. Author [Ano12, Ano13, Ano14, Ano16, Ano17]. automatic [HKS17]. Automorphism [AS18].

balance [Mil20]. Bar [Lon19]. Barry [ELPL^{+}18, Löw18, Löw16]. Base [dBSS16]. Base-complexity [dBSS16]. bases [AS18]. Behavior [SO13]. Bendixson [Joh20, Nor16]. between [FRS19, Fuj21, Mil20]. bi [BFRS21]. bi-embeddable [BFRS21]. block [Har15]. bound [HJMS16]. bounds [CKLZ20]. Bridging [EEH^{+}13].

Cantor [Joh20, Nor16]. Cardinal [GKT19]. case [GL17]. Categorical [MS13]. categoricity [AC17, BFRS21, CDHTM20]. ceers [AS19]. certain [Her20]. channels [ICDF17]. characterizations [HK18]. Characterizing [GL17]. choice [dBPS20]. circuits [BBF^{+}19]. classes [GKT19, Her20]. classification [Mil19]. classifications [dBSS16]. closed [Joh20]. coarse [HJMS16]. coding [DI17]. COH [CDHP20]. Coloring [FP17]. colouring [GJM^{+}19]. Combinatorial [DHH^{+}20, GL17]. communication [ICDF17]. Compactness [Kre12, Sha20]. compatible [GL17]. Complete [HHP14, Sha20]. Complexity [BDMP20, NNS14, BA15, FRS19, HK18, Koh19, MS17, Pat15, dBSS16]. component [GHM15]. Compositions [Goh20]. Compression [MT12]. Computability [AC17, BDKL19, HRW12, Löw16, Miy13, Sko13, TZ14, AHJ19, HJMS16]. Computable [CHM12, IP18, McN17, Lon19, MS19]. Computably [CDSS12, MS13]. computation [HK18, Mil20]. Computational [BDMP20, MS17, San18]. computations [BLPL18]. concerning [CDHP20]. cones [AHJ19]. Confluence [EEH^{+}13]. conjugacy [MS17]. Connected [Kih12, GHM15]. conservative [Tow15]. Constant [MT12]. constrained [GL17]. constructive [Fuj21, Nem20]. Content [BDMP20]. Continua [Kih12]. Continuity [Die13]. Continuous [TZ14]. Cooper [ELPL^{+}18, Löw16, Löw18]. copies [McN17]. copy [MS19]. Correspondence [STZDG13]. countable [Mil19, Sha20].

Dagstuhl [BDMP20]. Data [EEH^{+}13, PZ18]. Decidability [BA15]. definability [LP18]. definable [Mil17]. Degree [DGT13, Car16, MS19]. Degrees [BFRS21, CHS14, CDHTM20, FS14, Wri18, AS18, CD19, CLX18, DGH^{+}20, FT18, GS18, Her20, LP18, Sel20]. Dense [AHJ19, Nem20]. density [Ast15, HJMS16]. Derivative [HRW12]. Descartes [Mil20]. describing [Her20]. differentiable [Nem20]. differential [Mil20]. Dimension [ACDN17]. Diminishable [FFH^{+}20]. Discrete [TZ14].

Editorial [BBC^{+}12, BBBM16]. Effective [FvO16, HTKT19]. effectively [Joh20]. Efficient [Hit12]. elements [CDH^{+}19]. embeddable [BFRS21]. engine [Löw16]. Enumerable [CDSS12, AS18]. enumeration [GS18]. Equivalence [CHM12, FRS19]. equivalent [DHH^{+}20]. Ershov [CGH^{+}18].

Europe [Löw16]. evaluation [KN21]. Evaluations [STZDG13]. every [LL15]. exactly [CDH^{+}19]. existence [Fuj21, GHM15, Nem20]. expansions [Mil17]. Expressiveness [DM18]. extension [Mil20]. Extensions [DHS13, Tow15]. extraction [ACDN17].

Feasible [BBF^{+}19]. feasibly [KN21]. Fermat [MS13]. Field [DHS13]. Fields [MS13]. Finite [HW12, BH18]. Finitely [JKS18]. first [HT19]. first-order [HT19]. FO [Mil17]. forcing [Pat17]. Ford [BH18]. Forests [Her20]. forming [Ger20]. foundational [Mil20]. free [Sel20]. Fulkerson [BH18]. function [PZ18]. functional [San18]. Functions [FrKHNS14, Sko13, BBF^{+}19, BA15, Goh20, Har15, Her20, KN21, Nem20].

Gandy [San18]. Generalizations [BCNS14]. generated [JKS18]. Generator [AC12]. generic [Car17]. genericity [FT18]. Geometric [CHS14]. Gödel [Nor16]. Graph [EEH^{+}13, GHM15]. groups [JKS18].

halting [CD18, CD20]. hardness [GJM^{+}19]. Hierarchy [CJK^{+}13, CHM12, CGH^{+}18, GS18, SS15]. High [SO13]. higher [BS16]. higher-order [BS16]. Hindman [CKLZ20, CDH^{+}19]. Hyland [San18]. hyperimmunity [Pat17]. Hyperprojective [SS15].

II [Kri19]. immunity [Ast15, Pat18]. Implies [Die13]. Inapproximability [BCNS14]. incentive [GL17]. Incompleteness [Mor12]. Incomputability [Kih12]. Independence [STZDG13]. Index [Ano12, Ano13, Ano14, Ano16, Ano17]. induction [DHH^{+}20]. Infinite [Car17, Car16]. Information [HW12, STZDG13, DI17]. integer [Her16]. integer-valued [Her16]. integers [BA15]. intersection [IP18]. intuitionistic [BS16]. invariants [GKT19]. inversion [CLX18]. irrational [Kri17, Kri19]. isometry [MS19]. Isomorphism [CHS14, FS14, HHP14, Mil19, FT18]. issue [BDKL19, Löw18]. iterated [BA15]. iteration [Lon19]. Iterative [Pat17]. Ivan [Nik14].

Joins [AS19]. jump [Car17, CLX18, GS18].

Kalimullin [CLMS16]. Kernelization [EEH^{+}13, FFH^{+}20]. Kolmogorov [HK18, Hit12]. König [Die13].

Large [DGT13]. Last [MS13]. lattice [GMP19]. Layerwise [Miy13]. Leaf [Hir20]. Lemma [Die13]. limit [CDHTM20]. Limitations [Hit12]. limitwise [Har15]. linear [Sha20]. Linearisations [CGH^{+}18]. Lines [LL15]. Lipschitz [FrKHNS14]. Löf [BLPL18]. logic [BS16]. Lossiness [ICDF17]. Low [FS14, CD19]. Lowness [FT18, Her16, GKT19].

Machines [SS12, Car16, DM18]. management [Hir20]. Martin [BLPL18, HSY16]. Martin-Löf [BLPL18]. Mass [BS16]. Mathematics [DHS13, CDH^{+}19, FP17, Koh19, Pat15, Pat17, Pat18, San21]. maximal [Har15]. maximum [Tow15]. meanings [DM18]. Measures [STZDG13]. Measuring [BDMP20, FRS19]. mechanisms [GL17]. meets [AS19]. memory [Löw18]. minimal [AHJ19, Ger20]. missing [LL15]. modeled [ICDF17]. moduli [Koh19]. monadic [Spe15]. monotonicity [Har15]. Mortality [BA15]. multivalued [Goh20].

Natural [CHM12, DGT13, LP18]. Nets [San21]. network [BH18]. networks [PZ18]. Neumann [AC12]. Nikodym [HRW12]. non [GKT19]. non-lowness [GKT19]. nonstandard [San18]. Normalisation [AC12]. Notions [AZ19, DI17, Sko13]. nowhere [Nem20]. nowhere-differentiable [Nem20]. Number [AC12]. Numbers [CHM12, Kri17, Kri19]. Numerical [STZDG13].

Oberwolfach [BDKL19]. Objects [CHS14]. operations [FvO16]. operator [Car17]. Operators [TZ14, GMP19]. optimal [GL17]. oracles [Car17]. order [BS16, HT19, Spe15]. Orders [CDSS12, Pat18, Sha20]. Ordinal [SS12, HKS17]. ordinal-automatic [HKS17]. ordinals [CDHTM20, Wri18]. Overt [dBPS20].

pair [Ger20]. pairs [AHJ19, CLMS16]. Parameterized [BCNS14, FFH^{+}20]. Pareto [GL17]. part [Kri19]. Partial [CDSS12, Pat18]. PCAs [FvO16]. piecewise [BA15]. plain [HK18]. Planar [Kih12]. players [GL17]. point [LL15]. Pointed [BLPL18, HSY16]. points [IP18]. polynomial [FFH^{+}20]. Preface [BDKL19, Löw18]. principles [DHH^{+}20]. probabilistic [CD18]. problem [CD18, CDHP20, MS17, CD20]. problems [BS16, FFH^{+}20, Pat15]. problems1 [Muc16]. products [DGH^{+}20]. Projection [GMP19]. proof [Nem20]. pseudorandomness [ACDN17]. Pumping [HKS17].

qcb [SS15, dBSS16]. Quantum [AC12].

Radon [HRW12]. Ramsey [DGH^{+}20]. Random [AC12, Hit12, MT12, NNS14, HTKT19, LL15]. randomized [HK18]. Randomness [Car16, Mor12, Ast15, BLPL18, Her16]. rank [Nor16]. ranks [Joh20]. rationals [FP17]. REA [Ger20]. Real [Sko13]. reals [AZ19]. recognizability [Car17]. recognizable [Car17]. recursion [Joh20, Lon19]. Recursive [NNS14]. recursively [AS18]. Reducibility [CHM12, Hit12, Miy13, Fuj21, GKT19, Muc16]. Reduction [EEH^{+}13]. Reductions [HHP14, FRS19]. register [Car16]. regularity [Koh19]. Relations [CHM12, FRS19, Her20, Wri18]. Relative [Mor12]. representability [Kri17, Kri19]. Representations [KN21, SS12]. represented [Pau16]. restrictions [CKLZ20]. results [CDHP20, GJM^{+}19, Spe15]. Reverse [DHS13, CDH^{+}19, FP17, Koh19, Pat15, Pat17, Pat18, San21]. robust [DI17].

S. [ELPL^{+}18, Löw18]. satisfaction [Pat15]. satisfiability [Mil17]. search [LP18]. searchable [Nor16]. second [Spe15]. second-order [Spe15]. Selection [BCNS14]. Self [HW12]. Self-Information [HW12]. semiautomatic [JKS18]. semicomputability [AZ19]. Seminar [BDMP20]. Set [BCNS14, BBF^{+}19, Ger20, Mil17, Nor16]. Sets [HHP14, NNS14, Joh20]. Setting [CJK^{+}13]. Shared [Nik14]. Simply [Kih12]. small [BBF^{+}19]. Solovay [Miy13]. Some [CDHP20, Sko13, Spe15, CKLZ20]. Soskov [Nik14]. spaces [Pau16, SS15]. spaces1 [dBSS16]. special [BDKL19, Löw18]. Spectra [DGT13]. Splitting [CD19, CLX18]. Splittings [NNS14]. SRT22 [CDHP20]. statements [Fuj21]. statistical [CD20]. strategies [KN21]. Streams [TZ14, PZ18]. Strength [Kre12, CKLZ20, Sha20]. strict [FFH^{+}20]. Strings [Hit12]. Strong [HHP14, Muc16]. structure [AS19]. structures [AC17, HTKT19, HKS17, Mil19, Sel20]. subrecursive [Kri17, Kri19]. sufficiently [GL17]. sums [CDH^{+}19]. Surjective [GJM^{+}19]. symbolic [Mil20]. System [SO13].

Target [BCNS14]. Temperatures [SO13]. theorem [DGH^{+}20, HSY16, CKLZ20, CDH^{+}19, Die13, MS13]. theorems [CD19]. Theory [BDKL19, Car16, HT19, Joh20, Pau16]. thoughts [Nik14]. Tile [SO13]. Time [TZ14, Car16, Car17]. topological [AZ19, Her20, Pau16]. transducers [ICDF17]. Transfinite [BH18]. Transformation [EEH^{+}13]. Tree [SS12, HSY16]. Turing [CHS14, CLX18, LP18]. Two [GL17, CDH^{+}19]. type [FvO16, HT19]. Types [CHS14].

Ulm [HT19]. ultrahomogeneous [AC17]. Undecidability [Mil17]. Uniform [Die13]. uniqueness [Koh19]. upper [AHJ19]. using [HK18].

valued [Her16]. via [Lon19, MS13, SS12]. Volume [Ano12, Ano13, Ano14, Ano16, Ano17]. vs [CDHP20].

Wadge [Her20, Sel20]. Weak [Die13, Kre12, DHH^{+}20, Muc16]. weakly [AC17]. Weights [MT12]. Weihrauch [DGH^{+}20, Fuj21, GMP19, GKT19, Her20, Koh19]. words [Mil17]. Workshop [BDKL19].

References

Abbott:2012:NNQ

[AC12] Alastair A. Abbott and Cristian S. Calude. Von Neumann normalisation of a quantum random number generator. Computability, 1(1):59–83, ???? 2012. CODEN ???? ISSN 2211-3568 (print),

2211-3576 (electronic). URL http://iospress.metapress.com/
content/82271271358877v0.

Adams:2017:CCW

[AC17] Francis Adams and Douglas Cenzer. Computability and categoricity of weakly ultrahomogeneous structures. Computability, 6(4): 365–389, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Agrawal:2017:DPE

[ACDN17] Manindra Agrawal, Diptarka Chakraborty, Debarati Das, and Satyadev Nandakumar. Dimension, pseudorandomness and extraction of pseudorandomness. Computability, 6(3):277–305, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Astor:2019:DCU

[AHJ19] Eric P. Astor, Denis R. Hirschfeldt, and Carl G. Jockusch Jr. Dense computability, upper cones, and minimal pairs. Computability, 8 (2):155–177, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Anonymous:2012:AIV

[Ano12] Anonymous. Author index volume 1 (2012). Computability, 1(2): 181, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Anonymous:2013:AIV

[Ano13] Anonymous. Author index volume 2 (2013). Computability, 2 (2):141–142, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Anonymous:2014:AIV

[Ano14] Anonymous. Author index volume 3 (2014). Computability, 3(2): 147, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Anonymous:2016:AIV

[Ano16] Anonymous. Author index volume 5 (2016). Computability, 5(2): 197, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Anonymous:2017:AIV

[Ano17] Anonymous. Author index volume 6 (2017). Computability, 6(2): 209, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Ambos-Spies:2018:ABR

[AS18] Klaus Ambos-Spies. Automorphism bases for the recursively enumerable degrees. Computability, 7(2–3):237–258, ???? 2018. CO-DEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Andrews:2019:JMS

[AS19] Uri Andrews and Andea Sorbi. Joins and meets in the structure of ceers. Computability, 8(3–4):193–241, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Astor:2015:ADI

[Ast15] Eric P. Astor. Asymptotic density, immunity and randomness. Computability, 4(2):141–158, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Armstrong:2019:NST

[AZ19] Mark Armstrong and Jeffery Zucker. Notions of semicomputability in topological algebras over the reals. Computability, 8(1):1–26, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Ben-Amram:2015:MIP

[BA15] Amir M. Ben-Amram. Mortality of iterated piecewise affine functions over the integers: Decidability and complexity. Computability, 4(1):19–56, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Brattka:2016:E

[BBBM16] Vasco Brattka, Paola Bonizzoni, Olivier Bournez, and Elvira Mayordomo. Editorial. Computability, 5(1):1, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Brattka:2012:E

[BBC^{+}12] Vasco Brattka, Paola Bonizzoni, S. Barry Cooper, Benedikt Löwe, and Elvira Mayordomo. Editorial. Computability, 1(1):1–2, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Beckmann:2019:FSF

[BBF^{+}19] Arnold Beckmann, Sam Buss, Sy-David Friedman, Moritz Müller, and Neil Thapen. Feasible set functions have small circuits. Computability, 8(1):67–98, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Bazgan:2014:PIT

[BCNS14] Cristina Bazgan, Morgan Chopin, André Nichterlein, and Florian Sikora. Parameterized inapproximability of target set selection and generalizations. Computability, 3(2):135–145, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Brattka:2019:PSI

[BDKL19] Vasco Brattka, Rod Downey, Julia F. Knight, and Steffen Lempp. Preface of the special issue for the Oberwolfach Workshop on Computability Theory 2018. Computability, 8(3–4):191, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Brattka:2020:DSM

[BDMP20] Vasco Brattka, Damir D. Dzhafarov, Alberto Marcone, and Arno Pauly. Dagstuhl Seminar on Measuring the Complexity of Computational Content 2018. Computability, 9(3–4):167–168, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Bazhenov:2021:DBE

[BFRS21] Nikolay Bazhenov, Ekaterina Fokina, Dino Rossegger, and Luca San Mauro. Degrees of bi-embeddable categoricity. Computability, 10(1):1–16, ???? 2021. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Backman:2018:TFF

[BH18] Spencer Backman and Tony Huynh. Transfinite Ford–Fulkerson on a finite network. Computability, 7(4):341–347, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Barmpalias:2018:PCM

[BLPL18] George Barmpalias, Andrew Lewis-Pye, and Angsheng Li. Pointed computations and Martin-Löf randomness. Computability, 7(2–3): 171–177, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Basu:2016:MPI

[BS16] Sankha S. Basu and Stephen G. Simpson. Mass problems and intuitionistic higher-order logic. Computability, 5(1):29–47, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Carl:2016:RDT

[Car16] Merlin Carl. Randomness and degree theory for infinite time register machines. Computability, 5(2):181–196, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Carl:2017:ITR

[Car17] Merlin Carl. Infinite time recognizability from generic oracles and the recognizable jump operator. Computability, 6(3):223–247, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Calude:2018:PAA

[CD18] Cristian S. Calude and Monica Dumitrescu. A probabilistic anytime algorithm for the halting problem. Computability, 7(2–3): 259–271, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Chih:2019:STL

[CD19] Ellen Chih and Rod G. Downey. Splitting theorems and low degrees. Computability, 8(3–4):243–252, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Calude:2020:SAA

[CD20] Cristian S. Calude and Monica Dumitrescu. A statistical anytime algorithm for the Halting Problem. Computability, 9(2):155–166, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Csima:2019:RMH

[CDH^{+}19] Barbara F. Csima, Damir D. Dzhafarov, Denis R. Hirschfeldt, Carl G. Jockusch, Jr., Reed Solomon, and Linda Brown Westrick. The reverse mathematics of Hindman’s Theorem for sums of exactly two elements. Computability, 8(3–4):253–263, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Cholak:2020:SRC

[CDHP20] Peter A. Cholak, Damir D. Dzhafarov, Denis R. Hirschfeldt, and Ludovic Patey. Some results concerning the SRT22 vs. COH problem. Computability, 9(3–4):193–217, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Csima:2020:DCA

[CDHTM20] Barbara F. Csima, Michael Deveau, Matthew Harrison-Trainor, and Mohammad Assem Mahmoud. Degrees of categoricity above limit ordinals. Computability, 9(2):127–137, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Cholak:2012:CEP

[CDSS12] Peter A. Cholak, Damir D. Dzhafarov, Noah Schweber, and Richard A. Shore. Computably enumerable partial orders. Computability, 1(2):99–107, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Cooper:2018:LEH

[CGH^{+}18] S. Barry Cooper, James Gay, Charles M. Harris, Kyung Il Lee, and Anthony Morphett. Linearisations and the Ershov hierarchy. Computability, 7(2–3):143–169, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Coskey:2012:HER

[CHM12] Samuel Coskey, Joel David Hamkins, and Russell Miller. The hierarchy of equivalence relations on the natural numbers under computable reducibility. Computability, 1(1):15–38, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Calvert:2014:TDI

[CHS14] Wesley Calvert, Valentina Harizanov, and Alexandra Shlapentokh. Turing degrees of isomorphism types of geometric objects. Computability, 3(2):105–134, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Carson:2013:AHS

[CJK^{+}13] Jacob Carson, Jesse Johnson, Julia Knight, Karen Lange, Charles McCoy, and John Wallbaum. The arithmetical hierarchy in the setting of ω_{1}. Computability, 2(2):93–105, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Carlucci:2020:NBS

[CKLZ20] Lorenzo Carlucci, Leszek Aleksander Kołodziejczyk, Francesco Lepore, and Konrad Zdanowski. New bounds on the strength of some restrictions of Hindman’s Theorem. Computability, 9(2):139– 153, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Cai:2016:KP

[CLMS16] Mingzhong Cai, Steffen Lempp, Joseph S. Miller, and Mariya I. Soskova. On Kalimullin pairs. Computability, 5(2):111–126, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Cooper:2018:SJI

[CLX18] S. Barry Cooper, Angsheng Li, and Mingji Xia. Splitting and jump inversion in the Turing degrees. Computability, 7(2–3):133– 142, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

deBrecht:2020:OC

[dBPS20] Matthew de Brecht, Arno Pauly, and Matthias Schröder. Overt choice. Computability, 9(3–4):169–191, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

deBrecht:2016:BCC

[dBSS16] Matthew de Brecht, Matthias Schröder, and Victor Selivanov. Base-complexity classifications of qcb_{0}-spaces1. Computability, 5 (1):75–102, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Dzhafarov:2020:RTP

[DGH^{+}20] Damir D. Dzhafarov, Jun Le Goh, Denis R. Hirschfeldt, Ludovic Patey, and Arno Pauly. Ramsey’s theorem and products in the Weihrauch degrees. Computability, 9(2):85–110, ???? 2020. CO-DEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Diamondstone:2013:NLD

[DGT13] David Diamondstone, Noam Greenberg, and Daniel Turetsky. Natural large degree spectra. Computability, 2(1):1–8, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Davis:2020:CPE

[DHH^{+}20] Caleb Davis, Denis R. Hirschfeldt, Jeffry Hirst, Jake Pardo, Arno Pauly, and Keita Yokoyama. Combinatorial principles equivalent to weak induction. Computability, 9(3–4):219–229, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Dorais:2013:RMA

[DHS13] François G. Dorais, Jeffry Hirst, and Paul Shafer. Reverse mathematics and algebraic field extensions. Computability, 2(2):75–92, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Dzhafarov:2017:NRI

[DI17] Damir D. Dzhafarov and Gregory Igusa. Notions of robust information coding. Computability, 6(2):105–124, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Diener:2013:WKL

[Die13] Hannes Diener. Weak König’s Lemma implies the Uniform Continuity Theorem. Computability, 2(1):9–13, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Davidson:2018:EMM

[DM18] Joe Davidson and Greg Michaelson. Expressiveness, meanings and machines. Computability, 7(4):367–394, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Ehrig:2013:CDR

[EEH^{+}13] Hartmut Ehrig, Claudia Ermel, Falk Hüffner, Rolf Niedermeier, and Olga Runge. Confluence in data reduction: Bridging graph transformation and kernelization. Computability, 2(1):31–49, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Elwes:2018:BC

[ELPL^{+}18] Richard Elwes, Andy Lewis-Pye, Benedikt Löwe, Dugald Macpherson, Dag Normann, Andrea Sorbi, Alexandra A. Soskova, Mariya I. Soskova, Peter van Emde Boas, and Stan Wainer. S. Barry Cooper (1943–2015). Computability, 7(2–3):103–131, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Fernau:2020:DPP

[FFH^{+}20] Henning Fernau, Till Fluschnik, Danny Hermelin, Andreas Krebs, Hendrik Molter, and Rolf Niedermeier. Diminishable parameterized problems and strict polynomial kernelization. Computability, 9(1):1–24, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Frittaion:2017:CRR

[FP17] Emanuele Frittaion and Ludovic Patey. Coloring the rationals in reverse mathematics. Computability, 6(4):319–331, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Freer:2014:AAL

[FrKHNS14] Cameron Freer, Bjø rn Kjos-Hanssen, André Nies, and Frank Stephan. Algorithmic aspects of Lipschitz functions. Computability, 3(1):45–61, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Fokina:2019:MCR

[FRS19] Ekaterina Fokina, Dino Rossegger, and Luca San Mauro. Measuring the complexity of reductions between equivalence relations. Computability, 8(3–4):265–280, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Franklin:2014:DLI

[FS14] Johanna N. Y. Franklin and Reed Solomon. Degrees that are low for isomorphism. Computability, 3(2):73–89, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Franklin:2018:LID

[FT18] Johanna N. Y. Franklin and Dan Turetsky. Lowness for isomorphism and degrees of genericity. Computability, 7(1):1–6, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Fujiwara:2021:WCR

[Fuj21] Makoto Fujiwara. Weihrauch and constructive reducibility between existence statements. Computability, 10(1):17–30, ???? 2021. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Faber:2016:EOT

[FvO16] Eric Faber and Jaap van Oosten. Effective operations of type 2 in PCAs. Computability, 5(2):127–146, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Gerdes:2020:RSF

[Ger20] Peter Gerdes. An ω-REA set forming a minimal pair with 0 ∼′. Computability, 9(1):37–50, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Gura:2015:ECC

[GHM15] Kirill Gura, Jeffry L. Hirst, and Carl Mummert. On the existence of a connected component of a graph. Computability, 4(2):103– 117, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Golovach:2019:SCN

[GJM^{+}19] Petr A. Golovach, Matthew Johnson, Barnaby Martin, Daniël Paulusma, and Anthony Stewart. Surjective H-colouring: New hardness results. Computability, 8(1):27–42, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Greenberg:2019:CIN

[GKT19] Noam Greenberg, Rutger Kuyper, and Dan Turetsky. Cardinal invariants, non-lowness classes, and Weihrauch reducibility. Computability, 8(3–4):305–346, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Gonen:2017:CIC

[GL17] Rica Gonen and Anat Lerner. Characterizing incentive compatible, Pareto optimal and sufficiently anonymous constrained combinatorial mechanisms — two players case. Computability, 6(2): 183–208, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Gherardi:2019:POW

[GMP19] Guido Gherardi, Alberto Marcone, and Arno Pauly. Projection operators in the Weihrauch lattice. Computability, 8(3–4):281– 304, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Goh:2020:CMF

[Goh20] Jun Le Goh. Compositions of multivalued functions. Computability, 9(3–4):231–247, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Ganchev:2018:JHE

[GS18] Hristo A. Ganchev and Mariya I. Soskova. The jump hierarchy in the enumeration degrees. Computability, 7(2–3):179–188, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Harris:2015:LMM

[Har15] Charles M. Harris. On limitwise monotonicity and maximal block functions. Computability, 4(2):119–139, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Herbert:2016:LIV

[Her16] Ian Herbert. Lowness for integer-valued randomness. Computability, 5(2):103–109, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Hertling:2020:FDW

[Her20] Peter Hertling. Forests describing Wadge degrees and topological Weihrauch degrees of certain classes of functions and relations. Computability, 9(3–4):249–307, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Harkins:2014:SRI

[HHP14] Ryan C. Harkins, John M. Hitchcock, and A. Pavan. Strong reductions and isomorphism of complete sets. Computability, 3(2): 91–104, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Hirst:2020:LM

[Hir20] Jeffry L. Hirst. Leaf management. Computability, 9(3–4):309–314, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Hitchcock:2012:LER

[Hit12] John M. Hitchcock. Limitations of efficient reducibility to the Kolmogorov random strings. Computability, 1(1):39–43, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Hirschfeldt:2016:ADC

[HJMS16] Denis R. Hirschfeldt, Carl G. Jockusch Jr., Timothy H. McNicholl, and Paul E. Schupp. Asymptotic density and the coarse computability bound. Computability, 5(1):13–27, ???? 2016. CO-DEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Hirahara:2018:CRC

[HK18] Shuichi Hirahara and Akitoshi Kawamura. On characterizations of randomized computation using plain Kolmogorov complexity. Computability, 7(1):45–56, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Huschenbett:2017:POA

[HKS17] Martin Huschenbett, Alexander Kartzow, and Philipp Schlicht. Pumping for ordinal-automatic structures ^{1}. Computability, 6(2): 125–164, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Hoyrup:2012:CRN

[HRW12] Mathieu Hoyrup, Cristóbal Rojas, and Klaus Weihrauch. Computability of the Radon–Nikodym derivative. Computability, 1(1): 3–13, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Holzl:2016:MPT

[HSY16] Rupert Hölzl, Frank Stephan, and Liang Yu. On Martin’s pointed tree theorem. Computability, 5(2):147–157, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Harrison-Trainor:2019:FOT

[HT19] Matthew Harrison-Trainor. A first-order theory of Ulm type. Computability, 8(3–4):347–358, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Harrison-Trainor:2019:EAA

[HTKT19] Matthew Harrison-Trainor, Bakh Khoussainov, and Daniel Turetsky. Effective aspects of algorithmically random structures. Computability, 8(3–4):359–375, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Hirschfeldt:2012:FSI

[HW12] Denis R. Hirschfeldt and Rebecca Weber. Finite self-information. Computability, 1(1):85–98, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Ibarra:2017:LCC

[ICDF17] Oscar H. Ibarra, Cewei Cui, Zhe Dang, and Thomas R. Fischer. Lossiness of communication channels modeled by transducers. Computability, 6(2):165–181, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Iljazovic:2018:CIP

[IP18] Zvonko Iljazović and Bojan Pazek. Computable intersection points. Computability, 7(1):57–99, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Jain:2018:FGS

[JKS18] Sanjay Jain, Bakhadyr Khoussainov, and Frank Stephan. Finitely generated semiautomatic groups. Computability, 7(2–3):273–287, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Johnston:2020:CBR

[Joh20] Reese Johnston. Cantor–Bendixson ranks of effectively closed sets in admissible recursion theory. Computability, 9(2):111–125, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Kihara:2012:ISC

[Kih12] Takayuki Kihara. Incomputability of simply connected planar continua. Computability, 1(2):131–152, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Konecny:2021:RES

[KN21] Michal Konecný and Eike Neumann. Representations and evaluation strategies for feasibly approximable functions. Computability, 10(1):63–89, ???? 2021. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Kohlenbach:2019:RMW

[Koh19] Ulrich Kohlenbach. On the reverse mathematics and Weihrauch complexity of moduli of regularity and uniqueness. Computability, 8(3–4):377–387, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Kreuzer:2012:SWC

[Kre12] Alexander P. Kreuzer. On the strength of weak compactness. Computability, 1(2):171–179, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Kristiansen:2017:SRI

[Kri17] Lars Kristiansen. On subrecursive representability of irrational numbers. Computability, 6(3):249–276, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Kristiansen:2019:SRI

[Kri19] Lars Kristiansen. On subrecursive representability of irrational numbers, part II. Computability, 8(1):43–65, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Lutz:2015:LME

[LL15] Jack H. Lutz and Neil Lutz. Lines missing every random point. Computability, 4(2):85–102, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Longley:2019:BRC

[Lon19] John Longley. Bar recursion is not computable via iteration. Computability, 8(2):119–153, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Lowe:2016:BCE

[Löw16] Benedikt Löwe. Barry Cooper (1943–2015): The engine of Computability in Europe. Computability, 5(1):3–11, ???? 2016. CO-DEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Lowe:2018:PSI

[Löw18] Benedikt Löwe. Preface to the special issue in memory of S. Barry Cooper (1943–2015). Computability, 7(2–3):101, ???? 2018. CO-DEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Lewis-Pye:2018:SND

[LP18] Andrew E. M. Lewis-Pye. The search for natural definability in the Turing degrees. Computability, 7(2–3):189–235, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

McNicholl:2017:CC

[McN17] Timothy H. McNicholl. Computable copies of ^{ℓ}p. Computability, 6(4):391–408, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Milchior:2017:USE

[Mil17] Arthur Milchior. Undecidability of satisfiability of expansions of FO[<] over words with a FO[+]-definable set. Computability, 6 (4):333–363, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Miller:2019:ICC

[Mil19] Russell Miller. Isomorphism and classification for countable structures. Computability, 8(2):99–117, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Milici:2020:DED

[Mil20] Pietro Milici. A differential extension of Descartes’ foundational approach: a new balance between symbolic and analog computation. Computability, 9(1):51–83, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Miyabe:2013:CLC

[Miy13] Kenshi Miyabe. L^{1}-computability, layerwise computability and Solovay reducibility. Computability, 2(1):15–29, ???? 2013. CO-DEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Morphett:2012:IAR

[Mor12] Anthony Morphett. Incompleteness, approximation and relative randomness. Computability, 1(2):109–130, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Miller:2013:CCF

[MS13] Russell Miller and Hans Schoutens. Computably categorical fields via Fermat’s Last Theorem. Computability, 2(1):51–65, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Miasnikov:2017:CCC

[MS17] Alexei Miasnikov and Paul Schupp. Computational complexity and the conjugacy problem. Computability, 6(4):307–318, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

McNicholl:2019:IDC

[MS19] Timothy H. McNicholl and Donald M. Stull. The isometry degree of a computable copy of ℓp. Computability, 8(2):179–189, ???? 2019. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Merkle:2012:CCR

[MT12] Wolfgang Merkle and Jason Teutsch. Constant compression and random weights. Computability, 1(2):153–169, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Muchnik:2016:SWR

[Muc16] Albert A. Muchnik. Strong and weak reducibility of algorithmic problems1. Computability, 5(1):49–59, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Nemoto:2020:CPD

[Nem20] Takako Nemoto. A constructive proof of the dense existence of nowhere-differentiable functions in C[0, 1]. Computability, 9(3–4): 315–326, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Nikolova:2014:STI

[Nik14] Stela Nikolova. Shared thoughts on Ivan–Soskov (1954–2013). Computability, 3(1):63–72, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Ng:2014:CRS

[NNS14] Keng Meng Ng, André Nies, and Frank Stephan. The complexity of recursive splittings of random sets. Computability, 3(1):1–8, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Normann:2016:CBR

[Nor16] Dag Normann. On the Cantor–Bendixson rank of a set that is searchable in Gödel’s T. Computability, 5(1):61–74, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Patey:2015:CSP

[Pat15] Ludovic Patey. The complexity of satisfaction problems in reverse mathematics. Computability, 4(1):69–84, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Patey:2017:IFH

[Pat17] Ludovic Patey. Iterative forcing and hyperimmunity in reverse mathematics. Computability, 6(3):209–221, ???? 2017. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Patey:2018:POI

[Pat18] Ludovic Patey. Partial orders and immunity in reverse mathematics. Computability, 7(4):323–339, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Pauly:2016:TAT

[Pau16] Arno Pauly. On the topological aspects of the theory of represented spaces. Computability, 5(2):159–180, ???? 2016. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Pocas:2018:ANF

[PZ18] Diogo Poças and Jeffery Zucker. Analog networks on function data streams. Computability, 7(4):301–322, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Sanders:2018:GHF

[San18] Sam Sanders. The Gandy–Hyland functional and a computational aspect of nonstandard analysis. Computability, 7(1):7–43, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Sanders:2021:NRM

[San21] Sam Sanders. Nets and reverse mathematics. Computability, 10 (1):31–62, ???? 2021. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Selivanov:2020:WDF

[Sel20] Victor Selivanov. Q-Wadge degrees as free structures. Computability, 9(3–4):327–341, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Shafer:2020:SCC

[Sha20] Paul Shafer. The strength of compactness for countable complete linear orders. Computability, 9(1):25–36, ???? 2020. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Skordev:2013:SCN

[Sko13] Dimiter Skordev. On some computability notions for real functions. Computability, 2(1):67–73, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Seki:2013:BTA

[SO13] Shinnosuke Seki and Yasushi Okuno. On the behavior of tile assembly system at high temperatures. Computability, 2(2):107–124, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Speranski:2015:SNR

[Spe15] Stanislav O. Speranski. Some new results in monadic second-order arithmetic. Computability, 4(2):159–174, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Schlicht:2012:TRO

[SS12] Philipp Schlicht and Benjamin Seyfferth. Tree representations via ordinal machines. Computability, 1(1):45–57, ???? 2012. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Schroder:2015:HHQ

[SS15] Matthias Schröder and Victor Selivanov. Hyperprojective hierarchy of qcb_{0}-spaces. Computability, 4(1):1–17, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Soler-Toscano:2013:CIN

[STZDG13] Fernando Soler-Toscano, Hector Zenil, Jean-Paul Delahaye, and Nicolas Gauvrit. Correspondence and independence of numerical evaluations of algorithmic information measures. Computability, 2 (2):125–140, ???? 2013. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Towsner:2015:MCE

[Tow15] Henry Towsner. On maximum conservative extensions. Computability, 4(1):57–68, ???? 2015. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Tucker:2014:COC

[TZ14] J. V. Tucker and J. I. Zucker. Computability of operators on continuous and discrete time streams. Computability, 3(1):9–44, ???? 2014. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).

Wright:2018:DRO

[Wri18] Matthew Wright. Degrees of relations on ordinals. Computability, 7(4):349–365, ???? 2018. CODEN ???? ISSN 2211-3568 (print), 2211-3576 (electronic).